# XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge

### v2

@@ 6.2 XTransfer performance
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines. We run 15 rounds for each setting by default throughout the evaluation. Figure 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
@@
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Table 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Figure 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2. We keep the other experimental setups unchanged. Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
6.3 Comprehensive comparison
Cross-modality FSL performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2). We compare XTransfer against a range of SOTA baselines shown in Table 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (Zhou et al., 2023). All methods on target datasets are evaluated under the same cross-modality FSL settings (see Appendix A), with MetaSense (Gong et al., 2019) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Table 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Table 4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. In 3-shot setting, our method achieves slightly lower accuracy than the oracle on HHAR, Gesture, and Emotion. This indicates that extremely low samples may not fully represent the test data distribution, leading to less stable feature alignment. Since the layers are reused from pre-trained sources, the performance also depends on their quality and semantic overlap with targets. Notably, the flexibility of XTransfer allows further gains in such low-shot cases by using additional source models with higher quality or closer relevance. In summary, the results show that XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Table 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
@@ 6.3 Comprehensive comparison
HHAR WESAD Gesture Writing Emotion ChestX
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
@@ 7 Conclusion
This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design for human sensing applications. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing.
@@ A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing (Chen et al., 2019). In meta-training stage, FSL requires a large-scale source dataset from the same modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution (Triantafillou et al., 2020).. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format (Song et al., 2023). The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing (Gong et al., 2019). LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense (Gong et al., 2019) is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022). In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transfer pre-trained source models across different modalities.
@@ A.2 Setups in preliminary study
As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022). In Figure 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
@@ A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure & Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL (Jiang et al., 2022) and Original methods. Figure 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
@@ B.1 Default reshaping
B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques (Shen et al., 2022; Frankle & Carbin, 2019) typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
@@ B.4 Observation for designing repair rate growth model
B.5 Dynamic search range mechanism
Based on the estimated repair rate, LWS control can initially use a small, fixed search range to efficiently filter out most layer candidates before initiating SRR processing. However, because the updates to $rate_{n}^{est}$ are limited, the estimation may be inaccurate, potentially affecting selection decisions (e.g., missing the optimal layer candidate). To mitigate this risk, we develop a dynamic search range mechanism that adjusts the search range to balance accuracy and efficiency. This mechanism is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range. Here, $\hat{u}$ denotes a signed unit vector. If $rate_{n-1}^{est}$ is less than $rate_{n-1}$, it indicates that the estimation might be inaccurate and $\hat{u}$ will be positive to increase the range in searching for the optimal layer. Otherwise, the range will be reduced to improve search efficiency.
B.6 Post fine-tuning
@@ C.3 Impact of PCA components
C.4 Removal performance
We next evaluate how channel removal enhances the average S-score of selected layers. We continue to use the same setup as above under 5-shot settings. We employ L2 norm-based (i.e., L2-based) model pruning (Frankle & Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Figure 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.5 Impact of search depth

### v3

@@ 6.2 XTransfer performance
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance each layer’s S-score with different MMC dimensionalities. We use HHAR and the standard pre-trained ResNet18 on miniImageNet under 5-shot setting. The layer channel removal is disabled. We test with 4 different levels of MMC dimensionality (i.e., number of layer channels) ranging from 64 to 512. We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines. We run 15 rounds for each setting by default throughout the evaluation. Figure 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
@@
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Table 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. We compare different configurations: ResNet18 (Single) and Multi-ResNet18 (Multi) without efficient search enabled, Multi-ResNet18 (Multi-Pre) with only pre-search check using top-1 layer selection enabled, and Multi-ResNet18 (Multi-Efficient) with both pre-search check and dynamic search range enabled. As shown in Figure 4(b), Multi-Pre achieves a lowest search time on average across all n-shot settings, but it leads to degraded accuracy by 3.36% and 7.94% on average, respectively, compared to Multi and Multi-Efficient. It also fails to reach oracle-level accuracy and presents the highest standard deviation on average. The results indicate that relying solely on the pre-search check introduces search stability issues due to inaccurate or drifted repair value estimation. In contrast, Multi-Efficient provides a strong balance between accuracy and efficiency. It reduces search time by 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. While Multi-Efficient takes a longer search time than both Single and Multi-Pre, it successfully achieves superior accuracy and search stability, highlighting the benefits of using multiple source models and dynamic search range mechanism. Notably, even the search space is expanded 5 times more compared to Single, Multi-Efficient requires only 2.1 times more search time, indicating strong scalability enabled by the efficient search. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process while improving accuracy and stability.
@@
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2. We keep the other experimental setups unchanged and include 2-shot settings. Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average in 3- to 10-shot settings, compared to that of 69.6%, 64.3%, and 67% achieved by using Text, Audio, Sensing sources, respectively. This suggests that the latent features learned from Image, trained on larger datasets with broader class semantics, provide higher discriminability than those learned from other modalities. Notably, Image achieves lower average accuracy in 2- and 3-shot settings, suggesting that extremely low target data may not fully represent the test data distribution, leading to less stable alignment when using Image source. In contrast, both Text and Sensing result in higher average accuracy and reach the oracle, indicating that performance varies with source model quality and semantic relevance to the target. These results also show that XTransfer can flexibly reuse diverse source models across modalities to improve ultra low-shot performance. Moreover, the output models using Text or Sensing sources require only 0.16G FLOPs on average, as 1D kernels require much less computation than 2D kernels. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation results see Appendix C.
6.3 Comprehensive comparison
Cross-modality FSL performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2). We compare XTransfer against a range of SOTA baselines shown in Table 1. We report two variants with different sources: Our-Single, using a ResNet18 pre-trained on miniImageNet, and Our-Multi, using Multi-ResNet18 with five corresponding image datasets. We additionally report GPT2-small fine-tuning for comparison, following the parameter setting in (Zhou et al., 2023). All methods on target datasets are evaluated under the same cross-modality FSL settings (Appendix A.1), where only few target data are fine-tuned with sources. MetaSense (Gong et al., 2019) serves as the oracle, trained on Conv5 with sufficient target sensor data (Table 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Table 4 and Appendix Table 12, both our variants outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. While our variants are slightly below the oracle on HHAR, Gesture, and Emotion under 3-shot settings, XTransfer can further improve by reusing source models with higher quality or closer semantic relevance, as shown in our source-impact study. Overall, XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge. More results see Appendix C.
@@ 7 Conclusion
This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design. XTransfer reuses the power of pre-trained models as free sources and requires only few sensor data to restructure models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing applications.
@@ A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing (Chen et al., 2019). In meta-training stage, FSL requires a large-scale source dataset from the same modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^1] The support set and query set are assumed to be drawn from the same data distribution (Triantafillou et al., 2020).. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format (Song et al., 2023). The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing (Gong et al., 2019). LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense (Gong et al., 2019) is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022). In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transferring pre-trained source models across different modalities, using only few labeled target data and without assuming any additional unlabeled target data.
@@ A.2 Setups in preliminary study
As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022). In Figure 2(b), we apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights. It uses Earth Mover’s Distance (EMD) to measure similarity between source models and targets. We test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2. It also assesses FSL difficulty by fine-tuning source models on targets under 5-shot settings with LOOCV enabled. Since source models usually involve a large number of classes (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities.
@@ A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure and Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL (Jiang et al., 2022) and Original methods. Figure 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
@@ B.1 Default reshaping
B.2 PCA-based layer channel removal
In the Removal stage, layer channel removal is designed to further enhance layer-wise performance while optimizing resource usage. This is achieved by removing unnecessary layer channels at each layer. Existing pruning techniques (Shen et al., 2022; Frankle and Carbin, 2019) typically rely on layer channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess layer channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based layer channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for layer channel removal. Our objective is to remove unnecessary layer channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all layer channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
@@ B.4 Observation for designing repair rate growth model
B.5 Post fine-tuning
@@ C.3 Impact of PCA components
C.4 Non-linear feature alignment performance
We essentially use PCA to construct a fixed, orthogonal anchor feature space in which alignment is measured and the anchor-based repair loss is computed at each layer across modalities. The proposed connector (i.e., repair generator) dynamically maps target-modality features into this anchor space, enabling the non-linear feature alignment. To examine the impact of using a linear PCA anchor space and to compare the alignment performance, we evaluate our repair loss with and without PCA against a Maximum Mean Discrepancy (MMD) loss baseline (Wang et al., 2023). To set up, we keep using the standard ResNet18 on miniImageNet (Table 3) as sources on HHAR as target (Table 2) in 3-, 5- and 10-shot settings with all configurations identical except for the alignment loss. As shown in Table 9, our repair loss with PCA anchor space achieves the highest accuracy of 70.5% and the lowest ATR of 1.69 on average across all n-shot settings. Compared to our repair loss without PCA, MMD loss performs a lower accuracy on average. This indicates that MMD loss is limited to aligning the latent feature distribution when the feature space is highly complex due to strong cross-modality discrepancies and only few target data. In addition, adversarial alignment methods such as DAPN (Zhao et al., 2021) and MDDA (Zhao et al., 2020) also face similar difficulties under these conditions, as shown in Tables 4 and 12. Moreover, the PCA anchor space provides a more stable basis for alignment and consistently enhances our repair loss performance. As a result, these findings further validate the effectiveness of our method in cross-modality FSL settings.
Alignment Method 3-shot 5-shot 10-shot
Accuracy (%) ATR Accuracy (%) ATR Accuracy (%) ATR
MMD Loss (Wang et al., 2023) 55.78 1.63 65.89 2.01 68.52 1.83
Repair Loss 65.11 2.33 69.33 2.45 72.07 2.29
PCA + Repair Loss 58.96 1.42 71.85 1.56 80.74 2.09
Table 9: Comparison of non-linear feature alignment methods across n-shot settings.
C.5 Layer-wise convergence comparison
To further examine layer-wise convergence across the above alignment methods, we observe the feature alignment process in the initial layer. To normalize heterogeneous losses, we follow the strategy in (Li et al., 2020) and compute the loss descent (LD) relative to the initial loss, enabling a fair comparison of convergence rates across different loss functions (i.e., how much each loss decreases from its starting point), defined as $LD(t)=(L(0)-L(t))\mathbin{/}L(0)\times 100\%$, where $t$ denotes the epoch index. All experimental settings remain the same as in the previous setup. As shown in Table 10, our repair loss with the PCA anchor achieves the largest loss descent of 28.87%, which is 1.79 and 1.87 times faster than the repair loss without PCA and the MMD loss, respectively. Interestingly, MMD loss presents a large initial drop but quickly plateaus, while our repair loss starts more gradually and accelerates over training. In contrast, the PCA-based repair maintains a steady and consistent descent across epochs, highlighting the benefits of aligning features in a fixed, orthogonal anchor space. These findings further support that our method provides both effective feature alignment and practical training efficiency, consistent with the efficiency study in Section C.10.
Epoch (#) #10 #20 #30 #40 #50 #60 #70 #80 #90 #100
MMD Loss (Wang et al., 2023) 6.46% 10.0% 11.63% 12.81% 13.5% 14.3% 14.66% 14.98% 15.27% 15.41%
Repair Loss 0.03% 1.67% 2.66% 4.93% 5.26% 7.16% 8.06% 10.4% 14.66% 16.14%
PCA + Repair Loss 5.12% 13.45% 16.54% 19.23% 20.55% 21.3% 25.96% 26.21% 27.6% 28.87%
Table 10: Comparison of layer-wise convergence rates across alignment methods.
C.6 Removal performance
We next evaluate how layer channel removal enhances the average S-score of selected layers. We continue to use the same setup as above under 5-shot settings. We employ L2 norm-based (i.e., L2-based) model pruning (Frankle and Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where layer channel removal is disabled (i.e., Non-Pruning). Figure 7(a) demonstrates that the proposed layer channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the layer channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed layer channel removal further improves layer performance by accurately removing unnecessary layer channels at each layer.
C.7 Impact of search depth